# Floer homotopy theory for monotone Lagrangians

### v1

@@ Introduction
cf. Defs. 1.18, 1.19 and Thm. 1.24. It also induces generalized (co)homology theories associated to $\mathcal{F}$, under certain finiteness assumptions, cf. Def. 1.26 and Thm. 1.27.
@@
We remark that the existence of a $U$-brane actually forces $N_{\mu}$ to be even (Rmk. 2.24), so in fact $\tau_{\leq N_{\mu}-3}MU=\tau_{\leq N_{\mu}-4}MU$ since complex bordism is trivial in odd degrees. At present, due to a lack of development in the area of pro-spectra, we cannot vouch for the independence of the pro-spectrum with respect to the Hamiltonian and almost-complex structure, partly due to the ambiguity of what such a notion of equivalence of pro-spectra really means, cf. Rmk. 1.39 for a proposed definition, and the state of the literature. Before we introduce our application, we mention a certain class of Steenrod operations that lift to Postnikov truncations of complex bordism, that will be used in the application.
@@
    In Def. 1.19, we define the Steenrod algebra $\mathcal{A}^{*}_{M}(R)$ over an $\mathbb{E}_{1}$-ring spectrum $R$ with $M$-coefficients. Then, we state Thm. A in Thms. 1.24, 1.27, and prove them in §1.5, by constructing a pro-spectrum à la Cohen-Jones-Segal. The Milnor $Q_{i}$ operations of Fact C are also reviewed in this section, cf. Ex. 1.21 and Prop. 1.23.
@@
    We then define the notion of a $U$-brane (Def. 2.21) for the Lagrangian $L\subset M$, which by our index theorem endows $\mathcal{F}(M,L)$ with a $U$-structure, inducing a Steenrod algebra action on classical Floer cohomology ${\rm HF}^{*}(M,L)$ over the ring spectrum $\tau_{\leq N_{\mu}-3}MU$ (Thm. 2.26), cf. Thm. B above, as well as generalized cohomology theories with finite homotopy-support (Def. 1.26, Thm. 2.27).
@@
    We end with applications of our theory to $\mathbb{RP}^{n}\subset\mathbb{CP}^{n}$, showing how the Steenrod algebra structure provides further restrictions on the topology of possible clean intersections of $\mathbb{RP}^{n}$ with a Hamiltonian isotopy, cf. the preliminary results Thms. 2.44, 2.45 using only classical Floer cohomology, and then Thms. 2.47, 2.52, Cor. 2.48 using the Milnor $Q_{i}$ operations to strengthen the conclusion of Thm. 2.45. Finally, in Rmk. 2.53, we discuss further ramifications of these ideas.
@@ Acknowledgments
I would like to thank my advisor Ciprian Manolescu, for directing me and providing many useful suggestions throughout the writing process, as well as Mohammed Abouzaid, Kenneth Blakey, Andrew Blumberg, Sebastian Haney, Eleny Ionel, Bariş Kartal, Noah Porcelli, Ivan Smith, and Hiro Lee Tanaka, for illuminating comments and discussions. I am also greatly indebted to Stanford University, and to my sponsors for the William R. and Sara Hart Kimball Fellowship.
@@ 1.4. Steenrod action and generalized Floer (co)homology
Note bene 1.20 -- Terminology clash.
@@ 1.5. Construction thereof from a pro-spectrum
In this subsection, we prove the main results stated previously, namely Thms. 1.24, 1.27, via the Cohen-Jones-Segal (CJS) construction [13], [9, §5,§7]. There are two main difficulties which must be addressed, going beyond what the author wrote previously in [9]:
@@
Next, we deal with the case that $N<\infty$, so that the Pontrjagyn-Thom $\mathbb{E}_{1}$-map $\mathbf{E}_{*,*}^{+}\to R_{*,*}$ constructed in Def. 1.16 is only partially defined, when $\mu(i)-\mu(j)<N$. As a result, the equivalence relation in the definition of ${\rm CJS}^{u}$ is only partially defined. It turns out to be a bad idea to only impose this partial equivalence relation (i.e. its transitive closure), e.g. since Prop. 1.33 will not hold any longer. Instead, one must fill in the higher ${\rm PT}_{i,j}$ via obstruction theory:
@@
The result Thm. 1.24 now follows. Lastly, we note how this story parallels to generalized (co)homology theories, thereby proving Thm. 1.27.
@@
The analogous result to Prop. 1.37 holds for generalized (co)homology theories $Z^{*},Z_{*}$, given the finiteness condition Def. 1.26 is satisfied.
@@ 2.5. Independence of choices
In [9, §6], we have explained how to construct continuation maps on the level of Cohen-Jones-Segal constructions, lifting the continuation maps on the chain level, in the case that both flow categories have finite indexing sets $I$ and $J$. Even though in our case the flow categories are potentially unbounded both below and above, this does not really present an issue. As pointed out in the proof of Thm. 1.24, one can truncate the flow category $\mathcal{F}$ by only looking at indices $i<u<j$, for sufficiently small $i$ and sufficiently large $j$, and the underlying (co)homology will be independent in any finite range.
@@
at least when $Z$ is a $\tau_{\leq N_{\mu}-3}MU$-module with finite homotopy-support (cf. Def. 1.26). The cohomology has a well-defined action of the Steenrod algebra $\mathcal{A}^{*}_{Z}(\tau_{\leq N_{\mu}-3}MU)$, and there are both homological and cohomological Atiyah-Hirzebruch spectral sequences converging strongly to the generalized (co)homology just defined. These are independent of the choices of Hamiltonian and almost-complex structure.
@@ 2.7. Sample application: Projective spaces
taking $x_{r+d}\leftrightarrow\theta y_{d}$ (where $\theta$ is the Thom class of $\tilde{\xi}$, and the $x$’s and $y$’s follow the notations of Thms. 2.42, 2.45.) Cf. again Fig. 1. So for $10\leq d\leq n-\max(1,r)-2^{i+1}+$ 1, both the domain and codomain of the $Q_{i}$ operation lie in the aforementioned range, and in particular under this correspondence one has
@@
    More generally, if $\mathbb{RP}^{n}\cap\Phi_{1}(\mathbb{RP}^{n})$ is a disjoint union of $k>0$ connected components $C_{1}\sqcup\cdots\sqcup C_{k}$, which moreover satisfy $\dim C_{1}+\cdots+\dim C_{k}\leq n-k+1$, then they must have the same $\mathbb{F}_{2}$-cohomology as the real projective spaces of the respective dimensions, and likewise similar results should hold at the level of $Q_{i}$ Steenrod operations and $q_{i}$ characteristic classes.
@@
    As alluded to in Rmk.2.43, it should morally not be that difficult to complete the calculation of the Steenrod squares, provided one interprets them correctly in terms of real Gromov-Witten theory. This would likely sharpen the ranges in which we can say something about $Q_{i}$ and $q_{i}$.
@@
    Note that the non-vanishing of $Q_{i}$ implies the non-vanishing of all the ${\rm Sq}^{2^{j}}$, for all $j\leq i$. Thus, it is likely that Blakey [8]’s ideas could be made to work in the monotone setup, potentially providing new lower bounds for the number of (possibly degenerate) intersection points with a Hamiltonian isotopy.

### v2

@@ Introduction
cf. Defs. 1.181.19 and Thm. 1.25. It also induces generalized (co)homology theories associated to $\mathcal{F}$, under certain finiteness assumptions, cf. Def. 1.27 and Thm. 1.28.
@@
We remark that the existence of a $U$-brane actually forces $N_{\mu}$ to be even (Rmk. 2.24), so in fact $\tau_{\leq N_{\mu}-3}MU=\tau_{\leq N_{\mu}-4}MU$ since complex bordism is trivial in odd degrees. At present, due to a lack of development in the area of pro-spectra, we cannot vouch for the independence of the pro-spectrum with respect to the Hamiltonian and almost-complex structure, partly due to the ambiguity of what such a notion of equivalence of pro-spectra really means, cf.  Rmk. 1.40 for a proposed definition, and the state of the literature. Before we introduce our application, we mention a certain class of Steenrod operations that lift to Postnikov truncations of complex bordism, that will be used in the application.
@@
    In Def. 1.19, we define the Steenrod algebra $\mathcal{A}^{*}_{M}(R)$ over an $\mathbb{E}_{1}$-ring spectrum $R$ with $M$-coefficients. Then, we state Thm. A in Thms. 1.251.28, and prove them in §1.5, by constructing a pro-spectrum à la Cohen-Jones-Segal. The Milnor $Q_{i}$ operations of Fact C are also reviewed in this section, cf.  Ex. 1.21 and Prop. 1.23.
@@
    We then define the notion of a $U$-brane (Def. 2.21) for the Lagrangian $L\subset M$, which by our index theorem endows $\mathcal{F}(M,L)$ with a $U$-structure, inducing a Steenrod algebra action on classical Floer cohomology ${\rm HF}^{*}(M,L)$ over the ring spectrum $\tau_{\leq N_{\mu}-3}MU$ (Thm. 2.26), cf.  Thm. B above, as well as generalized cohomology theories with finite homotopy-support (Def. 1.27, Thm. 2.27).
@@
    We end with applications of our theory to $\mathbb{RP}^{n}\subset\mathbb{CP}^{n}$, showing how the Steenrod algebra structure provides further restrictions on the topology of possible clean intersections of $\mathbb{RP}^{n}$ with a Hamiltonian isotopy, cf.  the preliminary results Thms. 2.44, 2.45 using only classical Floer cohomology, and then Thms. 2.472.52, Cor. 2.48 using the Milnor $Q_{i}$ operations to strengthen the conclusion of Thm. 2.45. Finally, in Rmk. 2.53, we discuss further ramifications of these ideas.
@@ Acknowledgments
I would like to thank my advisor Ciprian Manolescu, supported by the Simons Investigator Grant, for directing me and providing many useful suggestions throughout the writing process, as well as Mohammed Abouzaid, Kenneth Blakey, Andrew Blumberg, Sebastian Haney, Eleny Ionel, Bariş Kartal, Noah Porcelli, Ivan Smith, and Hiro Lee Tanaka, for illuminating comments and discussions. I am also greatly indebted to my sponsors for the William R. and Sara Hart Kimball Fellowship offered by Stanford University, and to the Simons Foundation for organizing many conferences that benefitted my research.
@@ 1.4. Steenrod action and generalized Floer (co)homology
Nota bene 1.20 -- Terminology clash.
@@
Remark 1.24.
We conjecture and are quite confident that the $Q_{i}$’s are the only ordinary cohomology operations that lift to (indecomposable) cohomology operations over Postnikov truncations of $MU$, but we do not treat this subject here, as it would be too great of a digression. There are also plenty of non-zero cohomology operations over $\tau_{\leq r}MU$ that die even in the next $\tau_{\leq r+2}MU$, but this is a great subject for another paper.
@@ 1.5. Construction thereof from a pro-spectrum
In this subsection, we prove the main results stated previously, namely Thms. 1.251.28, via the Cohen-Jones-Segal (CJS) construction [13], [9, §5,§7]. There are two main difficulties which must be addressed, going beyond what the author wrote previously in [9]:
@@
Next, we deal with the case that $N<\infty$, so that the Pontrjagyn-Thom $\mathbb{E}_{1}$-map $\mathbf{E}_{*,*}^{+}\to R_{*,*}$ constructed in Def. 1.16 is only partially defined, when $\mu(i)-\mu(j)<N$. As a result, the equivalence relation in the definition of ${\rm CJS}^{u}$ is only partially defined. It turns out to be a bad idea to only impose this partial equivalence relation (i.e.  its transitive closure), e.g. since Prop. 1.34 will not hold any longer. Instead, one must fill in the higher ${\rm PT}_{i,j}$ via obstruction theory:
@@
The result Thm. 1.25 now follows. Lastly, we note how this story parallels to generalized (co)homology theories, thereby proving Thm. 1.28.
@@
The analogous result to Prop. 1.38 holds for generalized (co)homology theories $Z^{*},Z_{*}$, given the finiteness condition Def. 1.27 is satisfied.
@@
Remark 1.41 -- Obstruction theory and Steenrod operations.
The obstructions described in Prop. 1.35 in particular give sufficient conditions for when one can upgrade the Cohen-Jones-Segal construction from a Postnikov truncation $\tau_{\leq r}R$ to the next $\tau_{\leq r+1}R$. It would be interesting to know to what extent this also provides a necessary condition, since a priori this obstruction is built from choosing a specific cell structure for the module spectrum, as well as a particular $R$-orientation for the next moduli spaces.
We take the liberty here to speculate a bit more on this subject. Another way to prove that a $\tau_{\leq r}R$-module $M$ is not induced from $\tau_{\leq r+1}R$ is to show that $M$ has some nonzero cohomology operation $\alpha$ over $\tau_{\leq r}R$, such that $\alpha$ is in the kernel of $\mathcal{A}^{*}(\tau_{\leq r}R)\to\mathcal{A}^{*}(\tau_{\leq r+1}R)$. One wonders if this kind of Steenrod algebraic obstruction is at all related to the obstructions we got in Prop. 1.35.
We also note that the obstructions of Prop. 1.35, in the case of monotone Lagrangians (to be discussed in §2 shortly), should express some kind of real (homotopy) Gromov-Witten invariant. Classically, the curvature phenomena one obtains for minimal Maslov number $N_{\mu}=2$ translate algebraically into $d^{2}=c\cdot{\rm id}$, where $c$ is a count of holomorphic disks with boundary on the Lagrangian, hence exhibiting the obstruction to making ${\rm CF}_{*}$ a chain complex. If our speculations above are in any sense correct, then the real Gromov-Witten theory of the Lagrangian would be somehow reflected in the Steenrod algebra action. It would also show that the Floer homotopy type is homotopically “interesting,” as it does not extend to the next Postnikov truncation.
@@ 2.5. Independence of choices
In [9, §6], we have explained how to construct continuation maps on the level of Cohen-Jones-Segal constructions, lifting the continuation maps on the chain level, in the case that both flow categories have finite indexing sets $I$ and $J$. Even though in our case the flow categories are potentially unbounded both below and above, this does not really present an issue. As pointed out in the proof of Thm. 1.25, one can truncate the flow category $\mathcal{F}$ by only looking at indices $i<u<j$, for sufficiently small $i$ and sufficiently large $j$, and the underlying (co)homology will be independent in any finite range.
@@
at least when $Z$ is a $\tau_{\leq N_{\mu}-3}MU$-module with finite homotopy-support (cf.  Def. 1.27). The cohomology has a well-defined action of the Steenrod algebra $\mathcal{A}^{*}_{Z}(\tau_{\leq N_{\mu}-3}MU)$, and there are both homological and cohomological Atiyah-Hirzebruch spectral sequences converging strongly to the generalized (co)homology just defined. These are independent of the choices of Hamiltonian and almost-complex structure.
@@ 2.7. Sample application: Projective spaces
taking $x_{r+d}\leftrightarrow\theta y_{d}$ (where $\theta$ is the Thom class of $\tilde{\xi}$, and the $x$’s and $y$’s follow the notations of Thms. 2.422.45.) Cf.  again Fig. 1. So for $10\leq d\leq n-\max(1,r)-2^{i+1}+$ 1, both the domain and codomain of the $Q_{i}$ operation lie in the aforementioned range, and in particular under this correspondence one has
@@
    More generally, if $\mathbb{RP}^{n}\cap\Phi_{1}(\mathbb{RP}^{n})$ is a disjoint union of $k>0$ connected components $C_{1}\sqcup\cdots\sqcup C_{k}$, which moreover satisfy $\dim C_{1}+\cdots+\dim C_{k}\leq n-k+1$, then the inequality is in fact sharp, they must have the same $\mathbb{F}_{2}$-cohomology as the real projective spaces of the respective dimensions, and likewise similar results should hold at the level of $Q_{i}$ Steenrod operations and $q_{i}$ characteristic classes.
@@
    As alluded to in Rmk.2.43, it should morally not be that difficult to complete the calculation of the Steenrod squares, provided one interprets them correctly in terms of real Gromov-Witten theory. This would likely sharpen the ranges in which we can say something about $Q_{i}$ and $q_{i}$. Cf.  also Rmk. 1.41 about real Gromov-Witten theory potentially being related to obstructions seen at the level of Steenrod algebra actions on Floer cohomology.
@@
    Note that the non-vanishing of $Q_{i}$ implies the non-vanishing of all the ${\rm Sq}^{2^{j}}$, for all $j\leq i$. Thus, it is likely that Blakey [8]’s ideas could be adapted to the monotone setup, potentially providing new lower bounds for the number of (possibly degenerate) intersection points with a Hamiltonian isotopy.